0 0
Add a comment Improve this question Transcribed image text
Answer #1


Hot temperature T = 800 K
Cold temeprature T ' = 300 K
Maximum efficiency of the heat engine e = 1-( T ' / T )
= 1 - (3/8)
= 0.625
80% of maximum possible efficiency is 0.625*(80/100) = 0.5...

in percentage the answer is 0.5*100 = 50%
Answer #2

Nth=1-300/800

=0.625

80 % of 0.625

=0.5

=50 % ans